# MEGA: Masked Generative Autoencoder for Human Mesh Recovery

### v3

@@ Abstract
Human Mesh Recovery (HMR) from a single RGB image is a highly ambiguous problem, as an infinite set of 3D interpretations can explain the 2D observation equally well. Nevertheless, most HMR methods overlook this issue and make a single prediction without accounting for this ambiguity. A few approaches generate a distribution of human meshes, enabling the sampling of multiple predictions; however, none of them is competitive with the latest single-output model when making a single prediction. This work proposes a new approach based on masked generative modeling. By tokenizing the human pose and shape, we formulate the HMR task as generating a sequence of discrete tokens conditioned on an input image. We introduce MEGA, a MaskEd Generative Autoencoder trained to recover human meshes from images and partial human mesh token sequences. Given an image, our flexible generation scheme allows us to predict a single human mesh in deterministic mode or to generate multiple human meshes in stochastic mode. Experiments on in-the-wild benchmarks show that MEGA achieves state-of-the-art performance in deterministic and stochastic modes, outperforming single-output and multi-output approaches. Code and trained models will be released upon acceptance.
@@ 2.1 Human mesh recovery
Single output HMR. Since the release of the HMR [37] model, most approaches for recovering human meshes from images have been regression-based, using neural networks to make predictions directly from the image. These regression-based HMR methods can be categorized into parametric and non-parametric approaches. Parametric methods aim to recover the parameters of the SMPL model [76, 41, 48, 49, 27, 22, 42, 36, 91, 39]. They typically produce realistic predictions; however, some works have argued that the SMPL model parameter space is not the most suitable for predicting human meshes [17, 43, 14], leading to the development of non-parametric approaches. Non-parametric methods predict the coordinates of 3D vertices without relying on a parametric model. While earlier approaches used graph convolutional neural network architectures inspired by the mesh topology [43, 50], recent non-parametric models predominantly employ Transformers [13, 51, 21, 24, 58]. Although non-parametric methods yield accurate results, they can sometimes produce non-anthropomorphic meshes, particularly when training data is scarce [24].
In this work, instead of predicting SMPL model parameters or 3D coordinates, we aim to recover sequences of indices that can be decoded into a human mesh using the Mesh-VQ-VAE from [24]. Thus, the closest methods to ours are TokenHMR [22] and VQ-HPS [24], which rely on a quantized representation of human poses and meshes, respectively. However, the three methods have very distinct ways to address HMR: TokenHMR [22] is a regression method that uses tokens as an intermediate representation; VQ-HPS [24] is a classification method; MEGA is a generation method, which collapses to a classification method in the deterministic mode. Aside from this paradigm difference, MEGA takes partial mesh token sequences as input, while others are fed images. This allows us to first learn an unconditional human mesh generation model only on motion capture data (see Sec. C) before conditioning the generation with images. Ablation studies (see Tab. 1) show that this pre-training without any image conditioning significantly enhances the performance of MEGA.
@@ 3.3 Training strategy
Self-supervised pre-training. MEGA is pre-trained in a self-supervised manner on tokenized human meshes using a strategy similar to vector quantized masked autoencoders [31, 4, 69, 68]. The pre-training task involves reconstructing randomly masked human mesh tokens from a set of visible tokens. A variable masking rate is used such that $M=\lfloor N\cos(\frac{\pi\tau}{2})\rfloor$ with $\tau$ uniformly sampled from [0,1[[0,1[[ 0 , 1 [. The variable masking rate is critical for allowing MEGA to generate meshes iteratively in stochastic mode, as each step of the generation process involves predicting all tokens given a variable number of visible tokens (see Sec. 3.4). The sole loss used for pre-training is a cross-entropy loss computed from the reconstruction of the masked tokens.
Masked generative modeling for HMR. To train MEGA to predict tokenized canonical meshes from images, we extend the pre-training strategy by conditioning the decoder with the image embeddings (see Sec. 3.2 and Fig. 2). The human mesh tokens follow the same masking rate schedule, while the image embeddings remain fully visible. The only supervision for predicting canonical meshes is a cross-entropy loss, as in the pre-training stage. Compared to prior works in HMR that rely on multiple losses (on 3D keypoints, the 2D reprojection, and the SMPL parameters), this approach is straightforward and does not require hyperparameters tuning. For predicting the rotation and camera parameters, we use the Euclidean distance on the rotation matrix corresponding to the predicted 6D representation and the L1 loss on the reprojection of 2D joints extracted from the predicted oriented mesh, using the predicted perspective camera parameters.
@@
Table 1: Evaluation in deterministic mode. We evaluate MEGA on the 3DPW and EMDB datasets and compare it to the SOTA methods using metrics defined in Sec. 4.1 given in mm. $\dagger$ stands for additionally using 2D training data, and $\mp$ for additionally using 2D data and BEDLAM [7]. Methods in italic below the row ”MEGA” indicate the results of the ablation study.
@@ 4.1 Experimental setup
4.2 Deterministic inference mode
Quantitative evaluation. In  Tab. 1, we evaluate MEGA on the HMR task using the deterministic inference mode defined in Sec. 3.4, and compare its performance to SOTA methods on 3DPW [81] and EMDB [38]. Results for SOTA methods are taken from the corresponding papers when available or computed using official implementations and model weights. To ensure fairness, we only compare methods trained on standard datasets (see Sec. 4.1), with the exception of TokenHMR [22], which uses additional data but is included in the comparison because it also tokenized human pose. Note that the results for FastMETRO [13] and Virtual Marker [58] differ from those in their original papers because we present results without finetuning on the 3DPW dataset. Similar to [24, 22, 27], we chose not to use the 3DPW training set in order to measure generalization. Furthermore, it allows us to use the same model to evaluate MEGA on the occlusion dataset 3DPW-OCC [92] that contains sequences from the 3DPW training set.
@@ 4.2 Deterministic inference mode
Results on an occlusion dataset. We quantitatively evaluate MEGA on the occlusion dataset 3DPW-OCC [92] in Tab. 2 to evaluate the robustness of our approach to occlusions. Despite not being tailored for HMR on occluded images, MEGA surpasses all other methods trained on the same data (see Sec. 4.1). This performance could stem from MEGA’s self-attention mechanism among mesh tokens. While visible parts can be predicted by leveraging image embeddings, occluded parts heavily rely on visible body parts for accurate inference.
@@ 5 Conclusion
In this work, we explored self-supervised learning and masked generative modeling on human meshes for HMR. We introduced MEGA, a masked generative autoencoder designed to generate human meshes as discrete token sequences. MEGA’s flexible architecture and generation scheme enables the generation of diverse and realistic meshes and supports both single and multi-output HMR. An extensive evaluation demonstrates significant improvement over SOTA in both these domains.
@@ D Interpreting diversity in predictions
In the stochastic mode, MEGA makes diverse predictions given a single image. After making multiple predictions given an image, we can compute the standard deviation of the position of each vertex and interpret this value as a measure of the uncertainty in predictions. Indeed, if all samples are similar, we can conclude that the model is “certain” about this prediction. When the results are very diverse given a single image, we can interpret that as high uncertainty.

### v4

@@ Abstract
Human Mesh Recovery (HMR) from a single RGB image is a highly ambiguous problem, as an infinite set of 3D interpretations can explain the 2D observation equally well. Nevertheless, most HMR methods overlook this issue and make a single prediction without accounting for this ambiguity. A few approaches generate a distribution of human meshes, enabling the sampling of multiple predictions; however, none of them is competitive with the latest single-output model when making a single prediction. This work proposes a new approach based on masked generative modeling. By tokenizing the human pose and shape, we formulate the HMR task as generating a sequence of discrete tokens conditioned on an input image. We introduce MEGA, a MaskEd Generative Autoencoder trained to recover human meshes from images and partial human mesh token sequences. Given an image, our flexible generation scheme allows us to predict a single human mesh in deterministic mode or to generate multiple human meshes in stochastic mode. Experiments on in-the-wild benchmarks show that MEGA achieves state-of-the-art performance in deterministic and stochastic modes, outperforming single-output and multi-output approaches. See the project page at https://g-fiche.github.io/research-pages/mega/.
Work done at CentraleSupélec before joining Naver Labs Europe. Work done at IRI (CSIC-UPC) before joining Amazon.
@@ 2.1 Human mesh recovery
Single output HMR. Since the release of the HMR [37] model, most approaches for recovering human meshes from images have been regression-based, using neural networks to make predictions directly from the image. These regression-based HMR methods can be categorized into parametric and non-parametric approaches. Parametric methods aim to recover the parameters of the SMPL model [79, 42, 50, 52, 27, 22, 43, 36, 94, 39, 51]. They typically produce realistic predictions; however, some works have argued that the SMPL model parameter space is not the most suitable for predicting human meshes [17, 44, 14], leading to the development of non-parametric approaches. Non-parametric methods predict the coordinates of 3D vertices without relying on a parametric model. While earlier approaches used graph convolutional neural network architectures inspired by the mesh topology [44, 53], recent non-parametric models predominantly employ Transformers [13, 54, 21, 24, 61, 40]. Although non-parametric methods yield accurate results, they can sometimes produce non-anthropomorphic meshes, particularly when training data is scarce [24].
In this work, instead of predicting SMPL model parameters or 3D coordinates, we aim to recover sequences of token indices that can be decoded into a human mesh using the Mesh-VQ-VAE from [24]. Thus, the closest HMR method to ours is VQ-HPS [24], which also works with the token representation of the Mesh-VQ-VAE. However, there are also important differences between MEGA and VQ-HPS. First, VQ-HPS follows a deterministic classification-based mapping between images and mesh tokens. This contrasts with MEGA’s masked generative modeling approach, which allows multi-output predictions and unconditional human mesh generation (see Sec. C). Second, the architecture of MEGA differs from that of VQ-HPS. The two models’ encoders take different inputs: image patches for VQ-HPS and visible mesh tokens for MEGA. This key difference allows us to pre-train MEGA only with motion capture before conditioning the generation with images and to discard the encoder in the deterministic mode for faster inference. Ablation studies (see Tab. 1) show that this pre-training without any image conditioning significantly enhances the performance of MEGA. Another related HMR method is TokenHMR [22], which is a regression method that uses human pose tokens as an intermediate representation.
@@ 3.3 Training strategy
Self-supervised pre-training. MEGA is pre-trained in a self-supervised manner on tokenized human meshes using a strategy similar to vector quantized masked autoencoders [31, 4, 72, 71]. The pre-training task involves reconstructing randomly masked human mesh tokens from a set of visible tokens. A variable masking rate is used such that $M=\lfloor N\cos(\frac{\pi\tau}{2})\rfloor$ with $\tau$ uniformly sampled from [0,1[[0,1[[ 0 , 1 [. The variable masking rate is critical for allowing MEGA to generate meshes iteratively in stochastic mode, as each step of the generation process involves predicting all tokens given a variable number of visible tokens (see Sec. 3.4). The sole loss used for pre-training is a cross-entropy loss computed from the reconstruction of the masked tokens. At this stage of training, MEGA can be used for the unconditioned generation of random human meshes, as shown in Sec. C
Masked generative modeling for HMR. To train MEGA to predict tokenized canonical meshes from images, we extend the pre-training strategy by conditioning the decoder with the image embeddings (see Sec. 3.2 and Fig. 2). The human mesh tokens follow the same masking rate schedule, while the image embeddings remain fully visible. The only supervision for predicting canonical meshes is a cross-entropy loss, as in the pre-training stage. Compared to prior works in HMR that rely on multiple losses (on 3D keypoints, the 2D reprojection, and the SMPL parameters), this approach is straightforward and does not require hyperparameters tuning. Note that the performance of Mesh-VQ-VAE [24] is crucial because it provides the tokens we use as training targets. Hence, the reconstruction error is transferred to the ground truth used to train MEGA. Fortunately, the reconstruction error of Mesh-VQ-VAE is about one order of magnitude smaller than the estimation error of SOTA HMR models. For predicting the rotation and camera parameters, we use the Euclidean distance on the rotation matrix corresponding to the predicted 6D representation and the L1 loss on the reprojection of 2D joints extracted from the predicted oriented mesh, using the predicted perspective camera parameters.
@@
Table 1: Evaluation in deterministic mode. We evaluate MEGA on the 3DPW and EMDB datasets and compare it to the SOTA methods using metrics defined in Sec. 4.1 given in mm. $\dagger$ stands for additionally using 2D training data, and $\mp$ for additionally using 2D data and BEDLAM [7]. Methods in italic below the row "MEGA" indicate the results of the ablation study.
@@ 4.1 Experimental setup
4.2 Deterministic inference mode
Quantitative evaluation. In  Tab. 1, we evaluate MEGA on the HMR task using the deterministic inference mode defined in Sec. 3.4, and compare its performance to SOTA methods on 3DPW [84] and EMDB [38]. Results for SOTA methods are taken from the corresponding papers when available or computed using official implementations and model weights. To ensure fairness, we only compare methods trained on standard datasets (see Sec. 4.1), with the exception of TokenHMR [22], which uses additional data but is included in the comparison because it also tokenized human pose. Note that the results for FastMETRO [13] and Virtual Marker [61] differ from those in their original papers because we present results without finetuning on the 3DPW dataset. Similar to [24, 22, 27], we chose not to use the 3DPW training set in order to measure generalization. Furthermore, it allows us to use the same model to evaluate MEGA on the occlusion dataset 3DPW-OCC [95] (see below) that contains sequences from the 3DPW training set. Qualitative results of MEGA in the deterministic inference mode on in-the-wild datasets [84, 38] are shown in Sec. E.
Results on an occlusion dataset We quantitatively evaluate MEGA on the occlusion dataset 3DPW-OCC [95] in Tab. 2 to evaluate the robustness of our approach to occlusions. Despite not being tailored for HMR on occluded images, MEGA surpasses all other methods trained on the same data (see Sec. 4.1). This performance could stem from MEGA’s self-attention mechanism among mesh tokens. While visible parts can be predicted by leveraging image embeddings, occluded parts heavily rely on visible body parts for accurate inference.
@@ 5 Conclusion
In this work, we explored self-supervised learning and masked generative modeling on human meshes for HMR. We introduced MEGA, a masked generative autoencoder designed to generate human meshes as discrete token sequences. MEGA’s flexible architecture and generation scheme enables the generation of diverse and realistic meshes and supports both single and multi-output HMR. An extensive evaluation demonstrates significant improvement over SOTA in both these domains. Further discussions on the limitations and future works are available in Sec. F.
Acknowledgment
This study is part of the EUR DIGISPORT project supported by the ANR within the framework of the PIA France 2030 (ANR-18-EURE-0022). This work was performed using HPC resources from the “Mésocentre” computing center of CentraleSupélec, ENS Paris-Saclay, and Université Paris-Saclay supported by CNRS and Région Île-de-France. We would like to thank Samir Sadok for his valuable help with the implementation of the masked autoencoder.
@@ B Experimental details
Inference time In deterministic mode, with a batch size of 1, a forward pass takes 0.07 seconds with the HRNet version of MEGA on a GeForce GTX 1070 GPU. With similar settings, the ResNet version is real-time (0.03 seconds). In the stochastic mode, generating a single prediction with the ResNet model takes 0.04 seconds, which is still real-time, and generating 16 predictions takes 0.23 seconds.
@@ D Interpreting diversity in predictions
In the stochastic mode, MEGA makes diverse predictions given a single image. After making multiple predictions given an image, we can compute the standard deviation of the position of each vertex and interpret this value as a measure of the uncertainty in predictions. Indeed, if all samples are similar, we can conclude that the model is “certain" about this prediction. When the results are very diverse given a single image, we can interpret that as high uncertainty.
@@ E Qualitative results and failure cases
F Further discussions
Tokenizing vertices vs. SMPL parameters. Predicting the SMPL parameters presents weaknesses [24], such as error accumulation in the kinematic tree, that are addressed when working on the 3D vertices. The SMPL model parameters are attractive because they are a dense representation of human meshes, but once tokenized, this advantage is no longer leveraged. Even if the tokenization of Mesh-VQ-VAE starts from high-dimensional data, the obtained discrete representation is much denser than the representation proposed in TokenHMR in terms of sequence length ($54$ vs. $160$) and codebook size ($512\times 8$ vs. $2048\times 256$): it is easier to predict 54 tokens among 512 values than 160 tokens with 2048 possible values.
Potential applications. A direct application of MEGA is to generate solutions until we have a satisfactory answer, similar to LLMs. ScoreHypo [87] proposed a scoring network to select the best prediction among a range of outputs, increasing the accuracy of predictions. We can also choose the most suitable prediction depending on the use case: the solution that minimizes the re-projection error in sports applications for higher precision, the most visually appealing result in animation… The diversity of predictions can also be interpreted as a measure of per-vertex uncertainty [48] (see Sec. D).